# QVGen: Pushing the Limit of Quantized Video Generative Models

### v4

@@ Appendix I Comparison with Baselines for Huge DMs
33-bit quantization. As shown in Figs. A and B, our method QVGen far outperforms other baselines under $3$-bit quantization. Although $3$-bit quantization still introduces noticeable performance degradation, especially for huge DMs (see Figs. C and D), we believe QVGen represents a promising step toward practical ultra-low-bit video DMs.
@@
44-bit quantization. As depicted in Figs. E and F, previous QAT methods fail to deliver satisfactory results. In contrast, our method QVGen achieves video quality that closely approaches that of the full-precision model. Furthermore, for huge models (see Figs. G and H), QVGen consistently maintains high visual fidelity and effectively preserves generation quality.
Appendix M The Use of Large Language Models
@@
Figure A: Comparison of samples generated by full-precision and $3$-bit CogVideoX-22(Yang et al., 2025) models.
@@
Figure B: Comparison of samples generated by full-precision and $3$-bit Wan 1.31.3(WanTeam et al., 2025) models.
@@
Figure C: Comparison of samples generated by full-precision and $3$-bit CogVideoX1.5-55(Yang et al., 2025) models.
@@
Figure D: Comparison of samples generated by full-precision and $3$-bit Wan 1414(WanTeam et al., 2025) models.
@@
Figure E: Comparison of samples generated by full-precision and $4$-bit CogVideoX-22(Yang et al., 2025) models.
@@
Figure F: Comparison of samples generated by full-precision and $4$-bit Wan 1.31.3(WanTeam et al., 2025) models.
@@
Figure G: Comparison of samples generated by full-precision and $4$-bit CogVideoX1.5-55(Yang et al., 2025) models.
@@
Figure H: Comparison of samples generated by full-precision and $4$-bit Wan 1414(WanTeam et al., 2025) models.

### v5

@@ Appendix I Combination with SVDQuant
33-bit quantization. As shown in Figs. C and D, our method QVGen far outperforms other baselines under $3$-bit quantization. Although $3$-bit quantization still introduces noticeable performance degradation, especially for huge DMs (see Figs. E and F), we believe QVGen represents a promising step toward practical ultra-low-bit video DMs.
@@
44-bit quantization. As depicted in Figs. G and H, previous QAT methods fail to deliver satisfactory results. In contrast, our method QVGen achieves video quality that closely approaches that of the full-precision model. Furthermore, for huge models (see Figs. I and J), QVGen consistently maintains high visual fidelity and effectively preserves generation quality.
Appendix Q The Use of Large Language Models
@@
Figure C: Comparison of samples generated by full-precision and $3$-bit CogVideoX-22(Yang et al., 2025).
@@
Figure D: Comparison of samples generated by full-precision and $3$-bit Wan 1.31.3(WanTeam et al., 2025).
@@
Figure E: Comparison of samples generated by full-precision and $3$-bit CogVideoX1.5-55(Yang et al., 2025).
@@
Figure F: Comparison of samples generated by full-precision and $3$-bit Wan 1414(WanTeam et al., 2025).
@@
Figure G: Comparison of samples generated by full-precision and $4$-bit CogVideoX-22(Yang et al., 2025).
@@
Figure H: Comparison of samples generated by full-precision and $4$-bit Wan 1.31.3(WanTeam et al., 2025).
@@
Figure I: Comparison of samples generated by full-precision and $4$-bit CogVideoX1.5-55(Yang et al., 2025).
@@
Figure J: Comparison of samples generated by full-precision and $4$-bit Wan 1414(WanTeam et al., 2025).